# Jigsaw-Puzzles: From Seeing to Understanding to Reasoning in Vision-Language Models

### v1

@@ Abstract
Spatial reasoning is a core component of human cognition, enabling individuals to perceive, comprehend, and interact with the physical world. It relies on a nuanced understanding of spatial structures and inter-object relationships, serving as the foundation for complex reasoning and decision-making. To investigate whether current vision-language models (VLMs) exhibit similar capability, we introduce Jigsaw-Puzzles, a novel benchmark consisting of 1,100 carefully curated real-world images with high spatial complexity. Based on this dataset, we design five tasks to rigorously evaluate VLMs’ spatial perception, structural understanding, and reasoning capabilities, while deliberately minimizing reliance on domain-specific knowledge to better isolate and assess the general spatial reasoning capability. We conduct a comprehensive evaluation across 24 state-of-the-art VLMs. The results show that even the strongest model, Gemini-2.5-Pro, achieves only 77.14% overall accuracy and performs particularly poorly on the Order Generation task, with only 30.00% accuracy, far below the 90%+ performance achieved by human participants. This persistent gap underscores the need for continued progress, positioning Jigsaw-Puzzles as a challenging and diagnostic benchmark for advancing spatial reasoning research in VLMs.
@@
Zesen Lyu1,3, Dandan Zhang3, Wei Ye1, Fangdi Li2,3, Zhihang Jiang1,3, Yao Yang3thanks: Corresponding author. 1Hangzhou Institute for Advanced Study, UCAS 2Zhejiang University, 3Zhejiang Lab {lvzesen23, yewei23, jiangzhihang23}@mails.ucas.ac.cn, FdbendyLi@zju.edu.cn, {danae.zdd, yangyao}@zhejianglab.com
@@ 1 Introduction
However, existing benchmarks have yet to provide a comprehensive evaluation of spatial reasoning capability in VLMs under complex, real-world visual scenarios. Some works Fu et al. (2024); Li et al. (2024); Liu et al. (2024); Yue et al. (2024) focus primarily on foundational visual understanding by systematically evaluating perception, comprehension, and basic visual reasoning, revealing notable limitations in these areas. Although a few recent efforts Pothiraj et al. (2025); Stogiannidis et al. (2025); Ren et al. (2025); Tang et al. (2025) have attempted to evaluate the spatial reasoning capability of VLMs, they often rely on overly synthetic settings, task-specific constraints, or domain-dependent priors Song et al. (2025) (See Appendix A.1 for examples), limiting the capability to capture generalizable spatial reasoning under natural visual conditions. A truly effective evaluation of human-like spatial reasoning capability should model the task as a multi-stage cognitive process—beginning with perception, advancing through structural understanding, and culminating in high-level reasoning. Such reasoning must be grounded in the visual richness and ambiguity of real images, requiring the integration of spatial structural modeling and goal-directed reasoning Chen et al. (2024). Yet, this critical dimension of spatial cognition remains largely overlooked in existing benchmarks, underscoring the need for new benchmarks that move beyond narrow task formulations and embrace the full complexity of spatial reasoning.
To overcome the limitations of existing benchmarks in evaluating spatial reasoning capability of VLMs under real-world conditions, we introduce Jigsaw-Puzzles, a novel benchmark inspired by the cognitive mechanisms underlying human puzzle-solving. Puzzle-solving naturally reflects the multi-stage cognitive process Fissler et al. (2018), making it a compelling and effective testbed for spatial reasoning in VLMs.
@@
Capture Pothiraj et al. (2025) $\times$× $\times$× $\times$×
Mind the Gap Stogiannidis et al. (2025) $\times$× $\times$× $\times$×
VGRP Ren et al. (2025) $\times$× $\times$× $\times$×
LEGO-Puzzles Tang et al. (2025) $\times$×
@@ 3.2 Dataset Curation
Data Collection. We integrate data collection and quality control into a unified process. Starting from the CC3M Sharma et al. (2018) dataset, we apply task-specific filtering criteria—including minimum resolution and aspect ratio constraints—to construct an initial image pool of approximately 10,000 candidate images. Two human experts iteratively review the image pool while incrementally refining a shared set of filtering rules. Based on these evolving rules, they collaboratively filter the initial dataset to obtain the final set of high-quality, structurally diverse images. See Appendix A.2 for the rules pool. To enhance generalizability, we emphasize semantic and structural diversity throughout the dataset.
QA Generation. To support scalable and consistent QA pairs generation, each task type is associated with a specific construction template. QA pairs are automatically generated using these templates. Figure 3 illustrates simplified examples of the templates, full versions are provided in Appendix A.2.
@@ 4.1 Experimental Setting
Benchmark Models. We evaluate 24 VLMs on Jigsaw-Puzzles, covering a diverse range of model scales and training paradigms. For open-source models, we evaluate Qwen2-VL-72B Wang et al. (2024a), QvQ-72B-Preview Qwen (2024), Qwen2.5-VL-[7B/32B/72B] Bai et al. (2025), InternVL3-[8B/14B/38B/78B] Zhu et al. (2025), Kimi-VL-A3B-[Instruct/Thinking] Du et al. (2025), Phi-4-multimodal-instruct Abouelenin et al. (2025), Aya-Vision-[8B/32B] Dash et al. (2025), and Mistral-Small-3.1-24B-Instruct Mistral (2025). For proprietary models, we evaluate Claude-[3.5/3.7]-Sonnet Anthropic (2024), Gemini-[2.0/2.5]-Flash, Gemini-2.5-Flash-Thinking, Gemini-2.5-Pro Anil et al. (2023), GPT-4o, GPT-4o-mini Achiam et al. (2023), and Grok-2-Vision Grok (2024). Notably, QvQ-72B-Preview, Kimi-VL-A3B-Thinking, Gemini-2.5-Flash-Thinking, and Gemini-2.5-Pro are categorized as reasoning-enhanced models. All models, supporting multi-image input, are evaluated in a zero-shot setting with hardware scaled to their parameter size, see details in Appendix A.2.
@@
Models Easy Hard Easy Hard Connection Verification Anomaly Detection Order Restoration Overall
@@
Models Easy Hard Easy Hard Connection Verification Anomaly Detection Order Restoration Overall
@@ 5 Conclusion
Appendix A Appendix
A.1 Examples of Other Spatial Reasoning Benchmarks
@@ A.2 Dataset Curation
Hardware Setup for Evaluating VLMs. We evaluate open-source VLMs using hardware configurations scaled to model size. Models with fewer than 20B parameters run on a single NVIDIA A100 80GB GPU. Those between 20B and 40B use two NVIDIA A100 80GB GPUs, while models exceeding 40B are evaluated on four NVIDIA A100 80GB GPUs to meet their greater memory and computational demands.

### v4

@@ Abstract
Spatial reasoning is a core component of human cognition, enabling individuals to perceive, comprehend, and interact with the physical world. It relies on a nuanced understanding of spatial structures and inter-object relationships, serving as the foundation for complex reasoning and decision-making. To investigate whether current vision-language models (VLMs) exhibit similar capability, we introduce Jigsaw-Puzzles, a novel benchmark consisting of 1,100 carefully curated real-world images with high spatial complexity. Based on this dataset, we design five tasks to rigorously evaluate VLMs’ spatial perception, structural understanding, and reasoning capabilities, while deliberately minimizing reliance on domain-specific knowledge to better isolate and assess the general spatial reasoning capability. We conduct a comprehensive evaluation across 24 state-of-the-art VLMs. The results show that even the strongest model, Gemini-2.5-Pro, achieves only 77.14% overall accuracy and performs particularly poorly on the Order Generation task, with only 30.00% accuracy, far below the 90%+ performance achieved by human participants. This persistent gap underscores the need for continued progress, positioning Jigsaw-Puzzles as a challenging and diagnostic benchmark for advancing spatial reasoning research in VLMs. Our project page is at https://zesen01.github.io/jigsaw-puzzles.
@@
Zesen Lyu1,2, Dandan Zhang2, Wei Ye1, Fangdi Li2,3, Zhihang Jiang1,2, Yao Yang2thanks: Corresponding author. 1Hangzhou Institute for Advanced Study, UCAS 2Zhejiang Lab, 3Zhejiang University {lvzesen23, yewei23, jiangzhihang23}@mails.ucas.ac.cn, {danae.zdd, yangyao}@zhejianglab.com, FdbendyLi@zju.edu.cn
@@ 1 Introduction
However, existing benchmarks have yet to provide a comprehensive evaluation of spatial reasoning capability in VLMs under complex, real-world visual scenarios. Some works Fu et al. (2024); Li et al. (2024); Liu et al. (2024); Yue et al. (2024) focus primarily on foundational visual understanding by systematically evaluating perception, comprehension, and basic visual reasoning, revealing notable limitations in these areas. Although a few recent efforts Pothiraj et al. (2025); Stogiannidis et al. (2025); Ren et al. (2025); Tang et al. (2025) have attempted to evaluate the spatial reasoning capability of VLMs, they often rely on overly synthetic settings, task-specific constraints, or domain-dependent priors Song et al. (2025) (See Appendix A for examples), limiting the capability to capture generalizable spatial reasoning under natural visual conditions. A truly effective evaluation of human-like spatial reasoning capability should model the task as a multi-stage cognitive process—beginning with perception, advancing through structural understanding, and culminating in high-level reasoning. Such reasoning must be grounded in the visual richness and ambiguity of real images, requiring the integration of spatial structural modeling and goal-directed reasoning Chen et al. (2024). Yet, this critical dimension of spatial cognition remains largely overlooked in existing benchmarks, underscoring the need for new benchmarks that move beyond narrow task formulations and embrace the full complexity of spatial reasoning.
To overcome the limitations of existing benchmarks in evaluating spatial reasoning capability of VLMs under real-world conditions, we introduce Jigsaw-Puzzles, a novel benchmark inspired by the cognitive mechanisms underlying human puzzle-solving. Puzzle-solving naturally reflects the multi-stage cognitive process Fissler et al. (2018), making it a compelling testbed for spatial reasoning in VLMs. Unlike earlier studies that applied puzzle-solving primarily to visual representation learning Noroozi and Favaro (2016); Wang et al. (2022); Shen et al. (2024), our benchmark employs it to evaluate the spatial reasoning ability of VLMs under realistic visual conditions.
@@
Capture Pothiraj et al. (2025) $\times$ $\times$ $\times$
Mind the Gap Stogiannidis et al. (2025) $\times$ $\times$ $\times$
VGRP Ren et al. (2025) $\times$ $\times$ $\times$
LEGO-Puzzles Tang et al. (2025) $\times$
@@ 3.2 Dataset Curation
Data Collection. We integrate data collection and quality control into a unified process. Starting from the CC3M Sharma et al. (2018) dataset, we apply task-specific filtering criteria—including minimum resolution and aspect ratio constraints—to construct an initial image pool of approximately 10,000 candidate images. Two human experts iteratively review the image pool while incrementally refining a shared set of filtering rules. Based on these evolving rules, they collaboratively filter the initial dataset to obtain the final set of high-quality, structurally diverse images. See Appendix B for the rules pool. To enhance generalizability, we emphasize semantic and structural diversity throughout the dataset.
QA Generation. To support scalable and consistent QA pairs generation, each task type is associated with a specific construction template. QA pairs are automatically generated using these templates. Figure 3 illustrates simplified examples of the templates, full versions are provided in Appendix B.
@@ 4.1 Experimental Setting
Benchmark Models. We evaluate 24 VLMs on Jigsaw-Puzzles, covering a diverse range of model scales and training paradigms. For open-source models, we evaluate Qwen2-VL-72B Wang et al. (2024a), QvQ-72B-Preview Qwen (2024), Qwen2.5-VL-[7B/32B/72B] Bai et al. (2025), InternVL3-[8B/14B/38B/78B] Zhu et al. (2025), Kimi-VL-A3B-[Instruct/Thinking] Du et al. (2025), Phi-4-multimodal-instruct Abouelenin et al. (2025), Aya-Vision-[8B/32B] Dash et al. (2025), and Mistral-Small-3.1-24B-Instruct Mistral (2025). For proprietary models, we evaluate Claude-[3.5/3.7]-Sonnet Anthropic (2024), Gemini-[2.0/2.5]-Flash, Gemini-2.5-Flash-Thinking, Gemini-2.5-Pro Anil et al. (2023), GPT-4o, GPT-4o-mini Achiam et al. (2023), and Grok-2-Vision Grok (2024). Notably, QvQ-72B-Preview, Kimi-VL-A3B-Thinking, Gemini-2.5-Flash-Thinking, and Gemini-2.5-Pro are categorized as reasoning-enhanced models. All models, supporting multi-image input, are evaluated in a zero-shot setting with hardware scaled to their parameter size, see details in Appendix C.
@@
Models Easy Hard Easy Hard
Connection
Verification
Anomaly
Detection
Order
Restoration
Overall
@@
Models Easy Hard Easy Hard
Connection
Verification
Anomaly
Detection
Order
Restoration
Overall
@@ 5 Conclusion
Acknowledgments
The authors wish to thank the anonymous reviewers for their helpful comments. This work was supported by the Key R&D Program of Zhejiang (2024C01036).
Appendix A Examples of Other Spatial Reasoning Benchmarks
@@ Appendix B Dataset Curation
Appendix C Hardware Setup for Evaluating VLMs.
We evaluate open-source VLMs using hardware configurations scaled to model size. Models with fewer than 20B parameters run on a single NVIDIA A100 80GB GPU. Those between 20B and 40B use two NVIDIA A100 80GB GPUs, while models exceeding 40B are evaluated on four NVIDIA A100 80GB GPUs to meet their greater memory and computational demands.